# Fractionalization as an alternate to charge ordering in electronic insulators

### v1

@@ III.3 Construction of a fermionic SET of $G$
We can again turn to Cooper pairing to construct fermionic TOs, as illustrated in Fig. LABEL:fig:TOs and the previous section. The Cooper pairs are bosons at a filling $\nu/2=1/2q$ per unit cell. We then make a TO, $\mathcal{C}$, which is time reversal and translation invariant by considering the $\mathbb{Z}_{2q}$ gauge theory888stacked with $\{1,c\}$. of the Cooper pairs. This state is a fermionic Abelian TO with a GSD of $|\mathcal{C}|/2=4q^{2}$.
We can achieve a smaller quantum dimension if $q$ is odd, i.e. $q=2k+1$. Consider a band insulator which is made up of a state with one electron per site, and holes at a filling $\nu_{h}=1-\nu=2k/(2k+1)$. We then Cooper pair the holes to make a bosonic insulator at filling $\nu_{h}/2=k/(2k+1)$. A symmetry preserving bosonic insulating state can then be made with a $\mathbb{Z}_{2k+1}=\mathbb{Z}_{q}$ gauge theory of these paired holes. Thus for $q$ odd we have found a fermionic TO with GSD $|\mathcal{C}|/2=q^{2}$.
While the use of Cooper pairs to construct these states may not seem energetically natural if the dominant interaction is repulsive, in App. B we provide an alternate construction of these TOs using a parton construction which may yield a more natural route.
@@ III.4 Any minimal TO is Abelian
Having constructed a possible fermionic TO via Cooper pairing, we now prove that this is a minimal order consistent with $G$ in Eq. (4). We will do this by again following the strategy of Fig. 4. We begin by reviewing the basic Abelian anyons we know our theory posesses.
First, we have a background anyon $a$ with charge $Q_{a}=1/q$. As a consequence, the order $n_{a}$ must be an integer multiple of $2q$. Here the order $n_{a}$ is defined as the smallest non-zero integer such that $a^{n_{a}}=1$.
@@
Lastly, we know that $\theta_{v}=\pi\sigma_{H}\pmod{2\pi}$. Since our theory is time reversal invariant the Hall conductivity must be zero and thus $\theta_{v}=0$, i.e. $v$ is a charge-neutral boson.
@@ III.4.1 If $q$ is odd then the minimal order must be Abelian with a GSD of $q^{2}$
We start with $q$ odd as a warm-up. By our earlier arguments we know that the subset of Abelian anyons $\mathcal{A}\subseteq\mathcal{C}$ must contain at least three distinct anyons $a,v,c\in\mathcal{A}$. We consider all possible powers of these anyons in $\mathcal{A}$; the size of $|\mathcal{A}|$ is lower bounded by the number of their unique powers, as in Fig. 4. We will see that there are $2q^{2}$ Abelian anyons generated, proving that our earlier construction was minimal. Note again that we are counting the electron as an anyon.
@@ III.4.2 If $q$ is even then the minimal order must be Abelian with a GSD of $4q^{2}$
We will use a very similar argument as above, but we need to be a little more careful since the upper bound on our minimal number of anyons is now $8q^{2}$, where we are again counting the electron as an anyon for bookkeeping purposes. In the main text we first make the simplifying assumption that translation does not permute anyons unlike, e.g., Wen’s plaquette model Wen (2003). We show in App. D that under this assumption the time reversal invariance of our state will imply that $a=\mathcal{T}a$, and $a$ is thus a boson or fermion. We can then show that if $a=\mathcal{T}a$ there must be at least $8q^{2}$ Abelian anyons. In App. D we further allow for the possibility that translation permutes anyons and hence $a$ may not be the same as $\mathcal{T}a$, but nonetheless show that there must be at least $8q^{2}$ Abelian anyons.
We begin by noting that our logic for the case of $q$ odd revealed that there are $2q^{2}$ unique anyons in the set $\mathcal{S}=\{v^{m}a^{n}c^{k}\mid 0\leq m,n\leq q-1,k=0,1\}\subseteq\mathcal{C}$. Further, the only anyons in $\mathcal{S}$ which braid trivially with all other anyons in $\mathcal{S}$ are $1,c$. Now consider $a^{q}$. Since $a$ is a boson or fermion it must be the case that $a^{q}$ braids trivially with all powers of $a$. Moreover, since $a^{q}$ has integer charge, then it must braid trivially with all powers of $v$ as well. So then $a^{q}$ braids trivially with all anyons in $\mathcal{S}$. However, $a^{q}\neq 1$, because it has odd charge and the vacuum does not in a fermionic theory. Additionally $a^{q}\neq c$ since $a^{q}$ is a boson. So then we see $a^{q}\notin\mathcal{S}$.
@@
Next we know that since $a^{q}\neq 1,c$ but braids trivially with all anyons in $\mathcal{S}^{\prime}$ there must exist some $\gamma\in\mathcal{C}\setminus\mathcal{S}^{\prime}$ which braids non-trivially with $a^{q}$. Further, since $a^{2q}=1$, we must have that:
@@
Suppose then that $\gamma$ is non-Abelian. In App. D we prove that $\gamma$ can be taken to have the fusion products:
@@
We first sketch the argument for these fusion rules, with full proofs in App. D. Since $\gamma$ has $\pi$-braiding with $a^{q}$ we can always choose $v$ to be in the fusion product of $\gamma$ with itself. But since $\gamma$ is non-Abelian there must be some other anyon in this fusion product. By minimality the only possibilities are powers of $v,a,c$. Then the only other possibility by charge assignment is $va^{q}c$. This gives us the first fusion rule up to multiplicities of the fusion products, which we show in App. D must be one. The second fusion rule is a consequence of minimality as well; if it were not true there would be more than $8q^{2}$ anyons. Finally in App. D, we use a result in Lapa and Levin (2019) to show that time reversal invariance will require that $\gamma$ have $\theta_{\gamma}=0$.
@@
We can reach the conclusion that $\gamma$ must be Abelian through a different logic that does not directly rely on knowing that $\theta_{\gamma}=0$. Consider first the action of time reversal on $v$ and its powers. Clearly $v$ will transform to $v^{q-1}$. However for even $q$, $v^{\frac{q}{2}}$ will map to itself. We can then ask if it is a Kramers singlet or a Kramers doublet. Both possibilities are allowed and correspond to different SET states. First consider the case where it is a Kramers singlet. Then the state obtained by condensing $v$ will preserve time reversal. However as we have just seen, with a non-Abelian $\gamma$, this state will have the fusion rules of the Ising TO which is certainly not time reversal invariant, and we have a contradiction. Next consider the case where $v^{\frac{q}{2}}$ is a Kramers doublet. Now time reversal will have a complicated action on $v$, and the state obtained by condensing $v$ will not be time reversal invariant. Instead we first argue that for every TO where $v^{\frac{q}{2}}$ is a Kramers doublet, there is a partner TO where it is a Kramers singlet. To see this, we note that $v^{\frac{q}{2}}$ has $\pi$ braiding phase with $a$. If $a$ is a boson or a Kramers doublet fermion, we form a topological insulator of these particles. As is well-known Lu and Vishwanath (2012) this has the effect of toggling the $\pi$ flux seen by this particle between Kramers singlet and Kramers doublet. If instead $a$ is a Kramers singlet fermion, we repeat the same procedure with $ac$ which is a boson that has $\pi$ braiding with $v^{\frac{q}{2}}$. Importantly this procedure will not alter the quantum dimension of $\gamma$. Thus since we already argued that a phase with non-Abelian $\gamma$ and a Kramers singlet $v^{\frac{q}{2}}$ does not exist, it follows that such a phase also cannot exist with a Kramers doublet $v^{\frac{q}{2}}$.
@@ III.5 Uniqueness of the minimal TO
We have established that all minimal TOs consistent with the symmetry group $G$ of Eq. (4) must be Abelian and have an anyon count of $8q^{2}\ (2q^{2})$ for $q$ even (odd). As mentioned earlier this means that the definitions of minimality by either quantum dimension or ground state degeneracy will coincide, i.e. $mG_{\mathcal{D}}=mG_{\rm GSD}$. We will now show that our construction is always the unique minimal TO. We also comment on whether there is a unique symmetry enrichment of this TO. Throughout we will assume that translation does not permute anyons, so $a$ must be a boson or fermion. In App. E we present more general proofs without relying on this assumption.
@@ III.5.1 The case of $q$ odd
This case is very straightforward. We know that $a$ is a particle with charge $1/q$ and $a^{q}$ must braid trivially with everything. So then $a$ must be a fermion and $a^{q}=c$ by charge constraints. We further know that the vison is a charge zero boson and that all anyons in the minimal theory are powers of these two. Since the braiding of $a$ and $v$ is given by $\theta_{a,v}=2\pi/q$ the topological spin and braiding of all the anyons are then fully specified. It is simple to check that this will just be $\mathbb{Z}_{q}$ gauge theory.
In fact, this theory will not just be the unique minimal TO, but the unique minimal SET as well. To see this note that the charge assignment is totally fixed by the charges of $a$ and $v$. It is easy to see that the time reversal acts as $\mathcal{T}v=v^{-1},\mathcal{T}a=a$. Because $a^{q}=1$ and $q$ is odd, $a$ must be a Kramers singlet. Further, translation on $v$ is projective, and the corresponding phase is fully fixed by the mutual braiding between $a$ and $v$. Finally, we note that translation action on $a$ cannot be projective. For a projective translation we will have to place a background $v$ or some power of it at each plaquette center, as was considered in a time reversal breaking context in Song and Senthil (2022). However such a background is not time reversal invariant. Thus the translation action on $a$ (and its powers) is also totally fixed.
@@ III.5.2 The case of $q$ even
When $q$ is even we know that $a$ is an order $2q$ boson or fermion with charge $1/q$. Further, as discussed in the previous section there must exist some Abelian anyon $\gamma$ that has $\pi$ braiding with $a^{q}$ and all anyons in our theory must be made up of powers of $a,\gamma,c$. In order to fully specify the topological spin of our theory, we only need to know the topological spin of the Abelian $\gamma$. Without loss of generality we can assume that $\gamma^{2}=v$, so $\theta_{\gamma}=0,\pm\pi/2,\pi$ [^9] Note that earlier in Sec. III.4 we found $\theta_{\gamma}=0$ under the assumption that $\gamma$ was non-Abelian; having ruled out that scenario we now must determine $\theta_{\gamma}$ when $\gamma$ is Abelian., and $\theta_{\gamma,a}=\pi/q$. We can further take $\gamma$ to have charge zero; if it did not we could consider $\gamma a^{q}$ which would have the same braiding with $a$ and would also square to the vison.
We can now use a result from Lapa and Levin (2019) that
@@
for a $\mathrm{U}(1)_{f}$ symmetry enriched fermionic TO. Since the theory is time reversal invariant, both $c_{-}$ and $\sigma_{H}$ must vanish and the left-hand side will be equal to one. The right-hand side can then be evaluated using the fact that $04\theta_{\gamma}=$ 0 and $\theta_{a,\gamma}=\pi/q$. The end result is
@@
However, we note that in this case we have many choices of symmetry enrichment, as already alluded to in the previous section. Without loss of generality take the $e$ particle in $\mathbb{Z}_{2q}$ gauge to have charge $1/q$ and the $m$ particle to be charge neutral. Since $a$ is allowed to be a charge $1/q$ boson or fermion there are two possibilities: $a=e$, or $a=em^{q}$. The vison in this theory will always be given by $v=m^{2}$. In addition since we saw that the charge neutral $\gamma$ had to be a boson, then $\gamma=m$. It is also possible that $a$ or $m^{q}$, or both, may be a Kramers doublet since we have a Kramers singlet fermion [^10] In particular for Kramers singlet fermions with $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$ symmetry, the anomaly structure Wang and Senthil (2015) does not prevent this possibility.. Besides these possibilities we will not attempt to classify all possible SETs in this case, but will leave it to future works.
@@ IV Discussion
Here we explored the possiblity of a gapped quantum charge liquid (QCL) at a fractional filling of the underlying lattice which nevertheless does not break translation or time reversal symmetries. Such a QCL must display topological order (TO), and have excitations carrying fractional charge due to LSMOH constraints. Thus fractionalization can appear as an alternate to charge ordering at fractional lattice filling. Here we raised the question of the minimal topological order that is consistent with the symmetries and filling constraints, and established rigorous results. We showed that a fermionic QCL will generically be more complicated than the bosonic QCL at the same filling. A simple ilustration of our arguments is that a system of spinless fermions at $\nu=1/2$ cannot form a $\mathbb{Z}_{2}$ TO, but can form a $\mathbb{Z}_{4}$ TO. Thus the change in the statistics of the particles at filling $\nu=1/2$ must result in a dramatic shift in the TOs realized. Below we discuss a number of questions related to the physics of QCLs that naturally follow from our work.

### v2

@@ III.3 Construction of a fermionic SET of $G$
We can again turn to Cooper pairing to construct fermionic TOs, as illustrated in Fig. LABEL:fig:TOs and the previous section. The Cooper pairs are bosons at a filling $\nu/2=1/2q$ per unit cell. We then make a TO, $\mathcal{C}$, which is time-reversal and translation invariant by considering the $\mathbb{Z}_{2q}$ gauge theory [^13] Stacked with $\{1,c\}$. of the Cooper pairs. This state is a fermionic Abelian TO with a GSD of $|\mathcal{C}|/2=4q^{2}$.
We can achieve a smaller quantum dimension if $q$ is odd i.e., $q=2k+1$. Consider a band insulator which is made up of a state with one electron per site, and holes at a filling $\nu_{h}=1-\nu=2k/(2k+1)$. We then Cooper pair the holes to make a bosonic insulator at filling $\nu_{h}/2=k/(2k+1)$. A symmetry preserving bosonic insulating state can then be made with a $\mathbb{Z}_{2k+1}=\mathbb{Z}_{q}$ gauge theory of these paired holes. Thus for $q$ odd we have found a fermionic TO with GSD $|\mathcal{C}|/2=q^{2}$.
While the use of Cooper pairs to construct these states may not seem energetically natural if the dominant interaction is repulsive, in Appendix B we provide an alternate construction of these TOs using a parton construction which may yield a more natural route.
@@ III.4 Any minimal TO is Abelian
Having constructed a possible fermionic TO via Cooper pairing, we now prove that this is a minimal order consistent with $G$ in Eq. (4). We do this by again following the strategy of Fig. 4. We begin by reviewing the basic Abelian anyons we know our theory possesses.
First, we have a background anyon $a$ with charge $Q_{a}=1/q$. As a consequence, the order $n_{a}$ must be an integer multiple of $2q$. Here the order $n_{a}$ is defined as the smallest nonzero integer such that $a^{n_{a}}=1$.
@@
Lastly, we know that $\theta_{v}=\pi\sigma_{H}\pmod{2\pi}$. Since our theory is time-reversal invariant the Hall conductivity must be zero and thus $\theta_{v}=0$ i.e., $v$ is a charge-neutral boson.
@@ III.4.1 If $q$ is odd then the minimal order must be Abelian with a GSD of $q^{2}$
We start with $q$ odd as a warm-up. By our earlier arguments we know that the subset of Abelian anyons $\mathcal{A}\subseteq\mathcal{C}$ must contain at least three distinct anyons $a,v,c\in\mathcal{A}$. We consider all possible powers of these anyons in $\mathcal{A}$; the size of $|\mathcal{A}|$ is lower-bounded by the number of their unique powers, as in Fig. 4. We see that there are $2q^{2}$ Abelian anyons generated, proving that our earlier construction was minimal. Note again that we are counting the electron as an anyon.
@@ III.4.2 If $q$ is even then the minimal order must be Abelian with a GSD of $4q^{2}$
We use a very similar argument as above, but we need to be a little more careful since the upper bound on our minimal number of anyons is now $8q^{2}$, where we are again counting the electron as an anyon for bookkeeping purposes. In the main text we first make the simplifying assumption that translation does not permute anyons unlike, e.g., Wen’s plaquette model Wen (2003). We show in Appendix D that under this assumption the time-reversal invariance of our state will imply that $a=\mathcal{T}a$, and $a$ is thus a boson or fermion. We can then show that if $a=\mathcal{T}a$ there must be at least $8q^{2}$ Abelian anyons. In Appendix D we further allow for the possibility that translation permutes anyons and hence $a$ may not be the same as $\mathcal{T}a$, but nonetheless show that there must be at least $8q^{2}$ Abelian anyons.
We begin by noting that our logic for the case of $q$ odd revealed that there are $2q^{2}$ unique anyons in the set $\mathcal{S}=\{v^{m}a^{n}c^{k}\mid 0\leq m,n\leq q-1,k=0,1\}\subseteq\mathcal{C}$. Furthermore, the only anyons in $\mathcal{S}$ which braid trivially with all other anyons in $\mathcal{S}$ are $1,c$. Now consider $a^{q}$. Since $a$ is a boson or fermion it must be the case that $a^{q}$ braids trivially with all powers of $a$. Moreover, since $a^{q}$ has integer charge, then it must braid trivially with all powers of $v$ as well. So then $a^{q}$ braids trivially with all anyons in $\mathcal{S}$. However, $a^{q}\neq 1$, because it has odd charge and the vacuum does not in a fermionic theory. Additionally $a^{q}\neq c$ since $a^{q}$ is a boson. So then we see $a^{q}\notin\mathcal{S}$.
@@
Next we know that since $a^{q}\neq 1,c$ but braids trivially with all anyons in $\mathcal{S}^{\prime}$ there must exist some $\gamma\in\mathcal{C}\setminus\mathcal{S}^{\prime}$ which braids nontrivially with $a^{q}$. Furthermore, since $a^{2q}=1$, we must have that:
@@
Suppose then that $\gamma$ is non-Abelian. In Appendix D we prove that $\gamma$ can be taken to have the fusion products:
@@
We first sketch the argument for these fusion rules, with full proofs in Appendix D. Since $\gamma$ has $\pi$ braiding with $a^{q}$ we can always choose $v$ to be in the fusion product of $\gamma$ with itself. But since $\gamma$ is non-Abelian there must be some other anyon in this fusion product. By minimality the only possibilities are powers of $v,a,c$. Then the only other possibility by charge assignment is $va^{q}c$. This gives us the first fusion rule up to multiplicities of the fusion products, which we show in Appendix D must be one. The second fusion rule is a consequence of minimality as well; if it were not true there would be more than $8q^{2}$ anyons. Finally in Appendix D, we use a result in Ref. Lapa and Levin (2019) to show that time-reversal invariance will require that $\gamma$ have $\theta_{\gamma}=0$.
@@
We can reach the conclusion that $\gamma$ must be Abelian through a different logic that does not directly rely on knowing that $\theta_{\gamma}=0$. Consider first the action of time reversal on $v$ and its powers. Clearly $v$ will transform to $v^{q-1}$. However, for even $q$, $v^{\frac{q}{2}}$ will map to itself. We can then ask if it is a Kramers singlet or a Kramers doublet. Both possibilities are allowed and correspond to different SET states. First consider the case where it is a Kramers singlet. Then the state obtained by condensing $v$ will preserve time reversal. However, as we have just seen, with a non-Abelian $\gamma$, this state will have the fusion rules of the Ising TO which is certainly not time-reversal invariant, and we have a contradiction. Next consider the case where $v^{\frac{q}{2}}$ is a Kramers doublet. Now time reversal will have a complicated action on $v$, and the state obtained by condensing $v$ will not be time-reversal invariant. Instead we first argue that for every TO where $v^{\frac{q}{2}}$ is a Kramers doublet, there is a partner TO where it is a Kramers singlet. To see this, we note that $v^{\frac{q}{2}}$ has a $\pi$ braiding phase with $a$. If $a$ is a boson or a Kramers doublet fermion, we form a topological insulator of these particles. As is well known Lu and Vishwanath (2012), this has the effect of toggling the $\pi$ flux seen by this particle between Kramers singlet and Kramers doublet. If instead $a$ is a Kramers singlet fermion, we repeat the same procedure with $ac$ which is a boson that has $\pi$ braiding with $v^{\frac{q}{2}}$. Importantly, this procedure will not alter the quantum dimension of $\gamma$. Thus since we already argued that a phase with non-Abelian $\gamma$ and a Kramers singlet $v^{\frac{q}{2}}$ does not exist, it follows that such a phase also cannot exist with a Kramers doublet $v^{\frac{q}{2}}$.
@@ III.5 Uniqueness of the minimal TO
We have established that all minimal TOs consistent with the symmetry group $G$ of Eq. (4) must be Abelian and have an anyon count of $8q^{2}\ (2q^{2})$ for $q$ even (odd). As mentioned earlier this means that the definitions of minimality by either quantum dimension or ground-state degeneracy will coincide i.e., $mG^{(\nu)}_{\mathcal{D}}=mG^{(\nu)}_{\rm GSD}$. We now show that our construction is always the unique minimal TO. We also comment on whether there is a unique symmetry enrichment of this TO. Throughout we assume that translation does not permute anyons, so $a$ must be a boson or fermion. In Appendix E we present more general proofs without relying on this assumption.
@@ III.5.1 The case of $q$ odd
This case is very straightforward. We know that $a$ is a particle with charge $1/q$ and $a^{q}$ must braid trivially with everything. So then $a$ must be a fermion and $a^{q}=c$ by charge constraints. We further know that the vison is a charge-zero boson and that all anyons in the minimal theory are powers of these two. Since the braiding of $a$ and $v$ is given by $\theta_{a,v}=2\pi/q$ the topological spin and braiding of all the anyons are then fully specified. It is simple to check that this will just be $\mathbb{Z}_{q}$ gauge theory.
In fact, this theory will not just be the unique minimal TO, but the unique minimal SET as well. To see this note that the charge assignment is totally fixed by the charges of $a$ and $v$. It is easy to see that the time reversal acts as $\mathcal{T}v=v^{-1},\mathcal{T}a=a$. Because $a^{q}=1$ and $q$ is odd, $a$ must be a Kramers singlet. Furthermore, translation on $v$ is projective, and the corresponding phase is fully fixed by the mutual braiding between $a$ and $v$. Finally, we note that translation action on $a$ cannot be projective. For a projective translation we have to place a background $v$ or some power of it at each plaquette center, as was considered in a time reversal breaking context in Ref. Song and Senthil (2022). However such a background is not time-reversal invariant. Thus the translation action on $a$ (and its powers) is also totally fixed.
@@ III.5.2 The case of $q$ even
When $q$ is even we know that $a$ is an order $2q$ boson or fermion with charge $1/q$. Furthermore, as discussed in the previous section there must exist some Abelian anyon $\gamma$ that has $\pi$ braiding with $a^{q}$ and all anyons in our theory must be made up of powers of $a,\gamma,c$. To fully specify the topological spin of our theory, we only need to know the topological spin of the Abelian $\gamma$. Without loss of generality we can assume that $\gamma^{2}=v$, so $\theta_{\gamma}=0,\pm\pi/2,\pi$, [^14] Note that earlier in Sec. III.4 we found $\theta_{\gamma}=0$ under the assumption that $\gamma$ was non-Abelian; having ruled out that scenario we now must determine $\theta_{\gamma}$ when $\gamma$ is Abelian. and $\theta_{\gamma,a}=\pi/q$. We can further take $\gamma$ to have charge zero; if it did not we could consider $\gamma a^{q}$ which would have the same braiding with $a$ and would also square to the vison.
We can now use a result from Ref. Lapa and Levin (2019) that
@@
for a $\mathrm{U}(1)_{f}$ symmetry enriched fermionic TO. Since the theory is time-reversal invariant, both $c_{-}$ and $\sigma_{H}$ must vanish and the left-hand side will be equal to one. The right-hand side can then be evaluated using the fact that $04\theta_{\gamma}=$ 0 and $\theta_{a,\gamma}=\pi/q$. The end result is
@@
However, we note that in this case we have many choices of symmetry enrichment, as already alluded to in the previous section. Without loss of generality take the $e$ particle in $\mathbb{Z}_{2q}$ gauge to have charge $1/q$ and the $m$ particle to be charge neutral. Since $a$ is allowed to be a charge $1/q$ boson or fermion there are two possibilities: $a=e$, or $a=em^{q}$. The vison in this theory will always be given by $v=m^{2}$. In addition since we saw that the charge-neutral $\gamma$ had to be a boson, then $\gamma=m$. It is also possible that $a$ or $m^{q}$, or both, may be a Kramers doublet since we have a Kramers singlet fermion. [^15] In particular for Kramers singlet fermions with $\mathrm{U}(1)_{f}\rtimes\mathbb{Z}_{2}^{T}$ symmetry, the anomaly structure Wang and Senthil (2015) does not prevent this possibility. Besides these possibilities we will not attempt to classify all possible SETs in this case, but will leave it to future works.
@@ IV Discussion
Here we explored the possiblity of a gapped quantum charge liquid (QCL) at a fractional filling of the underlying lattice which nevertheless does not break translation or time-reversal symmetries. Such a QCL must display topological order (TO), and have excitations carrying fractional charge due to LSMOH constraints. Thus fractionalization can appear as an alternate to charge ordering at fractional lattice filling. Here we raised the question of the minimal topological order that is consistent with the symmetries and filling constraints, and established rigorous results. We showed that a fermionic QCL will generically be more complicated than the bosonic QCL at the same filling. A simple illustration of our arguments is that a system of spinless fermions at $\nu=1/2$ cannot form a $\mathbb{Z}_{2}$ TO, but can form a $\mathbb{Z}_{4}$ TO. Thus the change in the statistics of the particles at filling $\nu=1/2$ must result in a dramatic shift in the TOs realized. Below we discuss a number of questions related to the physics of QCLs that naturally follow from our work.